OpenStudy (anonymous):

please help \[x - \frac{ 20 }{ 3} = \frac{ 5 }{ 2 }x +\frac{ 5 }{ 6 }\]

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

You need ideas to get answers, right?

OpenStudy (anonymous):

linear equations

OpenStudy (skullpatrol):

What do you want to do?

OpenStudy (anonymous):

its problibly -51x-259=0 or x203=4(13x=14)

OpenStudy (anonymous):

how to do this problem with explanation

OpenStudy (anonymous):

\[x- \frac{20}{3} = \frac{5}{2}x+\frac{5}{6}\] \[ \rightarrow x- \frac{5}{2}x = +\frac{5}{6}+\frac{20}{3}\] \[ \rightarrow \frac{2x-5x}{2} = \frac{5}{6}+\frac{20 \times 2}{3 \times 2}\] \[ \rightarrow \frac{-3x}{2} = \frac{5}{6}+\frac{60}{6}\] \[ \rightarrow \frac{-3x}{2} = \frac{5+60}{6} \] \[ \rightarrow \frac{-3x}{2} = \frac{65}{6} \] \[ \rightarrow -3x = \frac{65 \times 2}{6} \] \[ \rightarrow -3x = \frac{65 }{3} \rightarrow x = \frac{65 }{3 \times (-3)} \[\rightarrow x = -\frac{65 }{9}\] @laal

OpenStudy (anonymous):

she has the right awnser laal

OpenStudy (anonymous):

thank you dpasingh very good explanation and very clean writing I got it

OpenStudy (anonymous):

yes, drewsheesley